OpenStudy (anonymous):

I NEED URGENT HELP ON THIS ONE (Question B in particular) A shower recess with dimensions 1.5m by 0.9m needs to have the back and two side walls tiled by Rudi to a height of 2 metres. (A) Calculate the area to be tiled in m^2? (The answer i came up with is 6.6m^2) (B) if tiles of dimensions 20cm by 20cm are to be used, exactly how many tiles will be required for the shower recess, including the ones that need to be cut. Disregard the grout.

OpenStudy (anonymous):

Well... You are given the base dimensions of 1.5m by 0.9m, and the hieght is 2m. So... Depending which is your back and which is your side walls, to find the area tiled, you would need to do XY+2*(ZY)=A X is your back length Y is your height Z is your side length A is your total area to be tiled

OpenStudy (anonymous):

Now once you find what your A is. you can take 20cm and convert this to m, which is 0.2m, so your tile area for 1 would be 0.2*0.2 = 0.04m, so once you find what your A is, if you do A/0.04=# Your # will be your total number of tiles used.

OpenStudy (anonymous):

what i did was 6.6 / 0.04 = 165 tiles but it has in the textbook 180 as the answer

OpenStudy (anonymous):

is that how im supposed to do it?

OpenStudy (anonymous):

Does it give a picture or anything? Depending what length you use as your back length and side length you will get different numbers, which did you use for which there?

OpenStudy (anonymous):

no there is no picture

OpenStudy (anonymous):

i used 1.5 x 2 to get the area of the back

OpenStudy (anonymous):

Well lets assume 1.5 is the sides and 0.9 is the back, so... 0.9*2+2*(1.5*2)=A=7.8m^2 now, 7.8/0.04= 195 tiles now if we switch them... 1.5*2+2*(0.9*2)=A=6.6m^2 now, 6.6/0.04= 165 tiles

OpenStudy (anonymous):

Does the book have the proper answer for part A?

OpenStudy (anonymous):

that is exactly what i did

OpenStudy (anonymous):

but in the textbook it says 180 tiles not 165

OpenStudy (anonymous):

yes part A is correct

OpenStudy (anonymous):

So it says part A is 6.6m^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there must be an error

OpenStudy (anonymous):

I mean if you made it work... Assuming: 180=6.6/X Then X is your tile area, which I get 0.03667, that makes no sense :P I would say the textbook has an error

OpenStudy (anonymous):

i have another question that is driving me nuts.

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

could you please help me put together a figure

OpenStudy (anonymous):

im really stuck on it

OpenStudy (anonymous):

What kind of figure?

OpenStudy (anonymous):

(a) Determine the exact total surface area of a sphere with radius square root of 2 metres (I know this one the answer is 8Pi square metre

OpenStudy (anonymous):

(b) An inverted cone with side length 4 metres is placed on top of the sphere so that the centre of its base is 0.5 metres above the centre of the sphere. Find the radius of the cone exactly?

OpenStudy (anonymous):

is the sphere drawn inside the inverted cone?